# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v7

@@ 4.3 Direct comparison
Table 1 summarises the principal distinctions among the three benchmarks. ARC and NeuroBench each presuppose the presence of high-level cognitive components—visual object recognition, manipulation and reasoning in the case of ARC, and comprehension and reasoning abilities in the case of NeuroBench. AGITB, by contrast, specifies elementary requirements that can be evaluated directly.
@@
Target reasoning type Human Universal Universal
Assumed cognitive priors Yes No No
Automation Limited Yes Yes
Teaching-to-the-test Yes No No
Intelligence level High Medium Low
@@
Spatial dimension $200\times 100$ 1 10
Temporal dimension Yes Yes Yes
Table 1: A list of core properties of the ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
@@ 5 Conclusion
Although AGITB is proposed as a general benchmark for artificial general intelligence, it is evidently grounded in empirical knowledge of neural function in the human brain. This grounding is justified by the fact that we have evidence that neuron-based systems—such as biological brains and artificial neural networks—are capable of producing general intelligence. In contrast, we have no demonstrated examples of AGI arising from fundamentally different architectures. Therefore, seeking more abstract or universal criteria may be premature, at least until the principles underlying natural, neuron-based intelligence are more fully understood.
Unlike conventional benchmarks that focus on high-level task performance—such as question answering or language translation—AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising twelve fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
AGITB requires models to begin from an unbiased initial state and to develop functionality solely through exposure to structured or unstructured input. This design reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity. In biological systems, high-level reasoning emerges not from symbolic manipulation but from the adaptive prediction of low-level sensory signals. This predictive process is more than pattern matching—it involves the gradual construction of signal-grounded knowledge that enables abstraction and generalisation over time.
AGITB is solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems. This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence. While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence. As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
// TODO While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture that can satisfy all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any non-AGI system, and it remains an open question whether such a solution is even theoretically possible.

### v8

@@ 4.3 Direct comparison
Table 1 highlights the key differences among the tasks used in the three benchmarks. Whereas ARC and NeuroBench presuppose or require models to exhibit high-level cognitive capacities (such as object recognition, spatial manipulation, and various forms of reasoning), AGITB instead focuses on minimal, precisely defined requirements that can be evaluated directly at the signal-processing level.
@@
AGI type Human Universal Universal
Cognitive priors Yes No No
Abstraction level High Medium Low
Formal AGI requirements No (qualitative) Partial Yes (explicit, testable)
@@
Input dimensionality $200\times 100$ (4 colours) 1 number 10 bits
Temporal sequence length 2 750+ 7+
Table 1: Core properties of ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.
@@ 5 Conclusion
Although AGITB is proposed as a general testbed and benchmark for artificial general intelligence, it is nonetheless grounded in empirical knowledge of neural function in the human brain. This grounding is motivated by the fact that neuron-based systems—most notably biological brains and, to a more limited extent, artificial neural networks—are the only systems known to support a broad range of flexible cognitive abilities. By contrast, there is currently no evidence that fundamentally different computational architectures can give rise to comparable forms of general intelligence. Pursuing more abstract or architecture-agnostic criteria may therefore be premature until the principles underlying natural, neuron-based intelligence are better understood.
Unlike conventional benchmarks that focus on high-level task performance, such as question answering or language translation, AGITB assesses whether a system exhibits behaviours thought to reflect core operational principles of the biological cortex. It focuses on low-level computational properties that are biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework comprising fourteen fundamental tests that evaluate a model’s ability to learn adaptively from raw input.
Unlike conventional benchmarks that target high-level task performance, such as question answering or language translation, AGITB evaluates whether a system exhibits behaviours associated with core operational principles of the biological cortex. Its focus is on low-level, biologically grounded computational properties that are believed to underlie the emergence of general intelligence. The testbed comprises fourteen tightly interdependent tests, each simple in isolation but collectively requiring the kind of adaptive, general-purpose learning expected of an AGI.
AGITB requires models to begin in an unbiased initial state and to acquire all functionality solely through exposure to structured or random input. This mirrors neuroscientific evidence that cortical learning is fundamentally input driven: neural circuits develop through experience, not through pre-encoded semantics. In biological systems, high-level cognition arises not from symbolic manipulation but from the continual adaptive prediction of low-level sensory signals. Such prediction is more than pattern matching; it supports the progressive construction of signal-grounded knowledge from which abstraction and generalisation can emerge.
AGITB is solvable by humans yet remains unsolved by classical algorithms and current state-of-the-art AI systems. This persistent performance gap provides empirical support for the claim that AGITB targets capabilities characteristic of general rather than narrow intelligence. Although the absence of an artificial solution does not constitute a formal proof of adequacy, the fact that humans succeed where machines do not indicates that the benchmark captures functionally relevant aspects of general intelligence. In this sense, AGITB serves as a discriminative test and a principled tool for steering the development of systems capable of genuinely general, adaptive learning.